OpenStudy (anonymous):

This a word problem about candy that I have gotten to two equations... 40x +100 y = 125 80x+84y=163 I dont know how to eliminate a variable or solve...thanks

OpenStudy (karatechopper):

Can you tell me what the Word problem was?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

40 candy bars and 100 drinks cost $125 and 80 candy bars and 84 drinks cost $163..how much is each candy bar?

OpenStudy (anonymous):

Solve for x or y in one of the equations

OpenStudy (anonymous):

x= 25 - 5/2 y ?

OpenStudy (anonymous):

x= 5/8 y?

OpenStudy (anonymous):

40x+100y=125 100y=-40x+125 y=-2/5x+5/4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now you know what y is so plug that into the other equation and solve for x 80x+84(-2/5x+5/4)=163

OpenStudy (anonymous):

Thx.

OpenStudy (anonymous):

right..Understand the theory-just seems overly complicated...usually the numbers are all nice and round and even... 84 times negative 2/5x + 5/4 makes my head spin..

OpenStudy (anonymous):

Well fractions just multiply straight across top and bottom....think of 84 as 84/1

OpenStudy (anonymous):

Thanks Caleb.

OpenStudy (anonymous):

What do you get for your answer?

OpenStudy (anonymous):

I dont...gonna let him finish the math and thanks for your help.

OpenStudy (anonymous):

75 cents and 1.25!!